OpenStudy (anonymous):

Find the value of x so that the line passing through the points will have the given slope: (x, 6) and (5, 3) with a slope of -3/4.

OpenStudy (aroub):

y=mx+b , m= -3/4 , ( 5,3) 3=-3/4(5)+b 3+15/4=b 12/4+15/4=b 27/4=b y=-3/4x+27/4